area of a parallelogram with vertices (13,18)( 43,18) (6,7)( 36,7)
Answer by LinnW(1048)

The top horizontal line is (13,18) between ( 43,18). Since 43-13 = 30,
the length of the base is 30.
Notice for the the pair (13,18) and (6,7) the height is 18-7 = 11.
So the base times the height is 30 * 11 = 330